# Electromagnetic energy calibration of the SoLid detector with horizontal muons

### v1

@@ 5.1.3 Results
The same fit strategy used for the simulation is applied to the calibration data obtained in the 2020 summer calibration campaign. An additional requirement based on the low multiplicity of the cubes (less than 5) is applied. Indeed, events with low multiplicity only accommodate the AC and 1 or 2 cubes for each annihilation gamma. This requirement therefore mitigates the contributions of proton recoils in the case of events with multiple neutrons. The best-fit result is shown in Figure 18. The value of the energy scale factor relative to the muon absolute calibration reads $0.9620\pm 0.0014$.
@@ 6 Crosscheck and validations
Note that the results reported in this section use both relative and absolute calibration performed with the horizontal muons. To begin with, one of the requirements posed to the calibration source was the statistical precision at the level of 1%. To reach it, the ROff data is gathered in 10-day periods. This sample size allows the precision of the method and the actual variations of the detector response to be aligned. Figure 19 shows an example of the statistical uncertainty distribution. The blue value corresponds to an uncertainty associated with one of the fibres in the main cube for each of the detection cells of the SoLid detector, whereas the orange distribution shows the same for one of the fibres impacted by the LL.
@@
Horizontal muons provide information on the time evolution of the detector response. The PVT is exposed to ageing, which decreases the performance of scintillating photon generation. Ageing is not homogeneous; hence, some of the fibres can be more exposed to it. Finally, the WLS fibres also undergo ageing, but the scale of the effect is much less significant, according to the manufacturer’s tolerances. All of these factors imply that the characteristics of light sharing evolve over time. To demonstrate it, we take one of the vertical fibres in the main cube of each detection cell. The fraction of light received by this fibre during the reference period is compared with the fraction received by the same fibre in the following periods. The plot on the left of Figure 20 shows this comparison for the reference period from the very beginning of data-taking in summer 2018 to periods 6, 12, and 18 months later. The mean value of the distribution that shifts towards the negative one is a clear measurement of the ageing of the detector. Simultaneously, the growing width of the distribution shows that ageing impacts the fibres differently; hence the importance of the frequent relative calibration of the detector. The same behaviour is obtained for the four main fibre distributions. The absolute energy scale can be used as an alternative approach to track the evolution of the detector response. The plot on the right of Figure 20 shows the evolution of the absolute energy scale factor. This is yet another demonstration of the ageing of PVT. The average effect is around 4-5% per year, which is slightly higher than expectation  [17]. It could be explained by a chiller failure that occurred in May 2019, when the temperature in the detector container increased to 50C instead of the usual 12C in standard operation. The incident is believed to have caused a degradation in the performance of the PVT. The final cross-check is done with a higher granularity. It explores the evolution of the response on the per-module level. The result is presented in Figure 21. The last detector module concentrates most of the dead fibres; thus, the average absolute energy scale in the module is expected to be lower, while it is determined with slightly worse precision. This effect is mitigated if the planes, placed after plane 45, are discarded. This is the region of the detector that accumulates most of the dead fibres. Without it, the performance of submodule 5 matches the performance of other modules.
@@ 6.1 Cosmogenic background determination
$\begin{gathered}\mu^{-}\;+\;^{12}C\;\longrightarrow\;^{12}B\;+\;\overline{\nu_{\mu}}\\ {}^{12}B\;\longrightarrow\;^{12}C\;+\;\overline{\nu_{e}}\;+\;e^{-}\;.\end{gathered}$ (6.1)
@@
12B is a well-known radioactive $\beta$ source. The half-life of the isotope is short (20.2 ms), and the end-point energy of subsequent $\beta$ decay is 13.4 MeV. Therefore, the energy distribution of these electrons completely covers the IBD energy range of interest. The selection of event candidates is based on the identification of muon tracks that stop inside the detector. An electromagnetic cluster is searched for, close in space and time to the reconstructed muon. To suppress the background, the energy of the cluster is required to be above 3 MeV. Furthermore, events with additional muon candidates within the 200 $\mu$s time window are removed. The yield of 12B is estimated from a fit of the time distance between the muon and the electromagnetic cluster. Finally, the energy spectrum reported in Figure 22 is determined after a statistical subtraction of the background using the sPlot [18] technique. The comparison between the measured spectrum and the reference from Geant4 shows a better agreement than 5% throughout the energy range. These results validate both relative and absolute calibration with horizontal muons. This standard candle background could not be observed in the SoLid data prior to the introduction of the muon calibration. The interested reader can find a comprehensive description of this analysis in [19].

### v2

@@ 5.1.3 Results
The same fit strategy used for the simulation is applied to the calibration data obtained in the 2020 summer calibration campaign. An additional requirement based on the low multiplicity of the cubes (less than 5) is applied (and by construction an event must have at least 3 cubes). Indeed, events with low multiplicity only accommodate the AC and 1 or 2 cubes for each annihilation gamma. This requirement therefore mitigates the contributions of proton recoils in the case of events with multiple neutrons. The best-fit result is shown in Figure 18. The value of the energy scale factor relative to the muon absolute calibration reads $0.9620\pm 0.0014$.
@@ 6 Crosscheck and validations
Note that the results reported in this section use both relative and absolute calibration performed with the horizontal muons. To begin with, one of the requirements of the calibration was a statistical precision at the level of 1%. To reach it, the ROff data are gathered in 10-day periods. This sample size allows the precision of the method and the actual variations of the detector response to be aligned. Figure 19 shows an example of the statistical uncertainty distribution. The blue value corresponds to an uncertainty associated with one of the fibres in the main cube for each of the detection cells of the SoLid detector, whereas the orange distribution shows the same for one of the fibres impacted by the LL.
@@
Horizontal muons provide information on the time evolution of the detector response. The PVT is exposed to ageing, which decreases the performance of scintillating photon generation. Ageing is not homogeneous; hence, some of the fibres can be more exposed to it. Finally, the WLS fibres also undergo ageing, but the scale of the effect is much less significant, according to the manufacturer’s information. All of these factors imply that the characteristics of light sharing evolve over time. To demonstrate it, we take one of the vertical fibres in the main cube of each detection cell. The fraction of light received by this fibre during the reference period is compared with the fraction received by the same fibre in the following periods. The plot on the left of Figure 20 shows this comparison for the reference period from the very beginning of data-taking in summer 2018 to periods 6, 12, and 18 months later. The mean value of the distribution that shifts towards the negative one is a clear measurement of the ageing of the detector. Simultaneously, the growing width of the distribution shows that ageing impacts the fibres differently; hence the importance of the frequent relative calibration of the detector. The same behaviour is obtained for the four main fibre distributions. The absolute energy scale can be used as an alternative approach to track the evolution of the detector response. The plot on the right of Figure 20 shows the evolution of the absolute energy scale factor. This is yet another demonstration of the ageing of PVT. The average effect is around 4-5% per year, which is slightly higher than expectation  [17]. It could be explained by a chiller failure that occurred in May 2019, when the temperature in the detector container increased to 50C instead of the usual 12C in standard operation. The incident is believed to have caused a degradation in the performance of the PVT. The final cross-check is done with a higher granularity. It explores the evolution of the response on the per-module level. The result is presented in Figure 21. The last detector module concentrates most of the dead fibres; thus, the average absolute energy scale in the module is expected to be lower, while it is determined with slightly worse precision. This effect is mitigated if planes beyond plane 45, where most of the dead fibres reside, are discarded. Without these planes, the performance of module 5 matches the performance of other modules.
@@ 6.1 Cosmogenic background determination
[…]
@@
12B is a well-known radioactive $\beta$ source. The half-life of the isotope is short (20.2 ms), and the end-point energy of subsequent $\beta$ decay is 13.4 MeV. Therefore, the energy distribution of these electrons completely covers the IBD energy range of interest. The selection of event candidates is based on the identification of muon tracks that stop inside the detector. Then an electromagnetic cluster, close in space and time to the reconstructed muon, is searched for. To suppress the background, the energy of the cluster is required to be above 3 MeV. Furthermore, events with additional muon candidates within the 200 $\mu$s time window are removed. It suppresses spallation neutrons that can mimic an electromagnetic cluster. The yield of 12B is estimated from a fit of the time distance between the muon and the electromagnetic cluster. Finally, the energy spectrum reported in Figure 22 is determined after a statistical subtraction of the background using the sPlot [18] technique. The comparison between the measured spectrum and the reference from Geant4 shows an agreement better than 5% throughout the energy range. These results validate both relative and absolute calibration with horizontal muons. This standard candle background could not be observed in the SoLid data prior to the introduction of the muon calibration. The interested reader can find a comprehensive description of this analysis in [19].